OpenStudy (anonymous):

A car dealer offers you two deals on a car that costs $16,000.Please calculate the monthly payment, given these two payment options the car dealer is offering. Payment Option 1: You can finance the car for 60 months with no interest if you make a $1,000 down payment. Payment Option 2: You can finance the car for 72 months (6 years) with 2% simple annual interest and no down payment. (Hint: To calculate simple annual interest, use the formula Interest = Principal * Rate * Time. Add the amount of interest to the price of the car.) Which monthly payment amount is lower? Please explain h

OpenStudy (anonymous):

Option 1: when you make down payment $1000 so the cost now is 16,000-1000=15,000 you divide for 60months for monthly payment=15,000/60=$250 Option2:this is amout for the car cost (2%*72months*16,000)=0.02*72*16,000=$23040 then you divide for 72 months for monthly cost=23040/72=$320 the Option 1=$250 the Option 2 =$320 so Option 1 is better